OpenStudy (anonymous):

From the local linearizations of e^x and sinx near x=0, write down the local linearization of the function e^x sins

ganeshie8 (ganeshie8):

Near 0, \(e^x \approx 1+x\) \(\sin x \approx x\)

ganeshie8 (ganeshie8):

Is your function : \(e^x \sin x\) like this ?

OpenStudy (anonymous):

yes

ganeshie8 (ganeshie8):

okay :) just multiply both the approximations and take only the linear terms : \(e^x \sin x \approx (1+x)x\) \(~~~~~~~~~~ \approx x + x^2\) \(~~~~~~~~~~ \approx x \)

ganeshie8 (ganeshie8):

see if that makes more or less sense...

OpenStudy (anonymous):

it does thanks!

ganeshie8 (ganeshie8):

np :)